The rate at which charge flows through a series circuit is inversely proportional to the amount of overall resistance in the circuit. Based on the results in Experiment 2, which one of the following rules accurately describes the effect of the number of bulbs upon overall current and overall resistance?

a. Increasing the number of bulbs will cause both the current and the resistance to increase. b. Increasing the number of bulbs will cause both the current and the resistance to decrease. c. Increasing the number of bulbs will cause the current to increase and the resistance to

decrease.

d. Increasing the number of bulbs will cause the current to decrease and the resistance to

increase.


1
Expert's answer
2021-12-15T11:15:41-0500

Let r "=" resistance of a bulb

Consider 3 bulbs in series with a battery of negligible internal resistance and of E.m.f. E


Resistance "=3r"

Current "=\frac{E}{3r}"


With 2 bulbs

Resistance "=2r"

Current "=\frac{E}{2r}"


With 1 bulb

Resistance "=r"

Current "=\frac{E}{r}"



As bulbs increases the current decreases and the resistance increases.